# “I Can See Forever!”: Evaluating Real-time VideoLLMs for Assisting Individuals with Visual Impairments

### v1

@@ Introduction
Based on our constructed benchmark dataset VisAssistDaily, we evaluate three models: GPT-4o, Zhipu, and VITA-1.5. The results are averaged from tests conducted by three individuals wearing blindfolds to simulate visual impairments (as shown Figure 2).
Table 1 presents the TSR across the three models. GPT-4o performs the best in most tasks, especially within the “Home Life Tasks” and “Basic Skills”, where it achieves 100% TSR on multiple tasks. Further, the gap between its TSR in English and Chinese is minimal, indicating balanced multilingual capabilities. Zhipu performs slightly below GPT-4o but still shows strong results in both task sets, achieving an average TSR of 55.22% and 73.48%, respectively. However, its performance varies significantly across tasks. For example, it records a low success rate of only 33.33% in “Road Walking (EN)” and just 16.67% in “Cane Techniques (EN)”, indicating challenges in handling certain task-specific scenarios. VITA-1.5 shows the weakest overall performance, with a TSR of only 53.30% in English and 61.18% in Chinese. The model demonstrates inconsistent behavior across languages for different tasks. For instance, it achieves 100% TSR in “Housework (CH)”, outperforming its 65.00% in “Housework (EN)”, but only 33.33% in “Reaching Destination (EN)”, reflecting unstable multilingual task performance.
Table 2 shows the models’ performance across various evaluation metrics. Overall, GPT-4o ranks the highest, achieving the highest TSR and LC, with the lowest ARL in English. It also records the lowest APC in CH tasks. Zhipu ranks second, with a solid average TSR of 73.48% under CH input and moderate ARL (2.66) and APC (3.52). However, it struggles to maintain consistent language output, often responding in Chinese to EN prompts. VITA-1.5 performs the worst overall across all models.
@@
In error case 1, GPT-4o demonstrates spatial reasoning limitations, mislocating a cane positioned on the left side of the room by describing it as being on the right and near a distant black chair. This reflects confusion between object co-occurrence and actual spatial proximity, as well as unreliable left-right orientation. In error case 2, Zhipu incorrectly localizes the user at the top of a staircase despite clear contextual cues indicating the bottom position. The model further described a flat hallway ahead, omitting the presence of stairs. This reflects a failure in spatial grounding and may result in safety risks in assistive navigation contexts.
[…]
@@
Participants. We recruit 8 volunteers through social media platforms, including RedNotes and Weibo, to participate in the user study in either an online or in-person manner. Each participant is compensated at an hourly rate of 15 dollars. As shown in Table 3, the participants are in a diverse range of ages (from 21 to 90). Among them, U1 to U5 are individuals who are severely visually impaired and are almost completely blind, while U6 to U8 are sighted individuals who wear blindfolds during study to simulate visual impairment.
@@
Task. The evaluation tasks are divided into two categories: closed-world tasks and open-world tasks (user-defined goals). Closed-world tasks include: 1. Finding a specific item in a room; 2. Identifying a book and helping to determine its title and table of contents; 3. Going upstairs, with the model guiding users by describing step height and assisting with navigation; 4. Entering a room and having the model describe the surroundings; 5. Walking outdoors, with the model providing feedback on road conditions. Open-world tasks include finding multiple similar objects, crossing the street, finding medicine bottles, judging road conditions and crossing traffic lights, finding tools in a workspace, assisting with learning, experiencing recreational activities, and expanding the scope of learning.
Performance Metrics. We use only TSR as the main metric, since some participants are not comfortable with recording video, making it difficult to collect other measurements. In addition, we use a Satisfaction score to capture overall user experience, which is rated by each participant after completing the tasks. The scoring criteria are detailed in Table 4.
@@
Performance Analysis. As shown in Table 3, we find that participants achieve high task success rates (TSR) in both closed-world and open-world settings. Notably, users U2, U6, U7, and U8 achieve a closed-world TSR of 100.00%. Besides, users U5, U4, and U3 achieve closed-world TSRs of 80.00%, 80.00%, and 75.00%, respectively. In the open-world settings, users U5, U6, and U7 achieve a TSR of 100.00%. Also, users U4, U8, and U1 attain TSRs of 83.00%, 80.00%, and 75.00%, respectively. These results suggest that Zhipu performs well in most scenarios, demonstrating high TSR across a broad range of tasks. However, a notable exception is observed with user U1 and U3, who have 60.00% and 50% closed-world TSR. This lower performance is likely due to Zhipu’s limitations in distinguishing between similar objects. Despite these exceptions, all users acknowledge the capabilities of Zhipu and are generally satisfied with the model, as demonstrated by their satisfaction scores are all above $6$. This indicates that, overall, users find the model capable of handling complex tasks to a reasonable competence.
Feedback from Participants. Participants’ feedback indicates that their needs align closely with daily living activities. These include meal preparation (U1), medication management (U1), personal hygiene (U1), mobility (U2, U3, U5), recreational activities (U5), and educational support (U4, U7). The main challenges they encounter when seeking technological assistance involve navigation tasks such as finding directions (U1), recognizing rooms and locations (U3, U5), crossing streets (U1, U2), and interpreting traffic signals (U2, U3). Participants also highlight the need for object localization, especially when locating items in the home, workplace, or leisure environments (U1, U2, U3, U4, U5, U6), along with assistance in conducting detailed environmental assessments and providing scene descriptions (U3, U4, U6, U8). Test results show that the assistive system performs inconsistently across different scenarios. Object and environmental descriptions are often inaccurate, particularly in crowded or complex settings (U4, U6, U8). Specific limitations include imprecise spatial descriptions (U4, U6), incorrect or incomplete object recognition (U4, U6, U8), lack of detail in descriptions needed by visually impaired users (U4, U8), difficulty converting visual information into actionable instructions (U4, U6), occasional incorrect confirmation responses (U4), and challenges with voice interaction, especially in noisy environments (U4, U8). Although some participants report high success in specific contexts such as locating daily objects (U1, U2, U3, U6), participating in recreational activities (U5), supporting educational tasks (U4, U7), and understanding environmental context (U8), user satisfaction varies. Concerns include the speed of interaction (U1, U8), difficulty understanding long or complex instructions (U7, U8), and the importance of continuous reminders related to visual impairment (U8). In addition, U4 participated in a longer in-depth interview. He pointed out two prominent issues encountered during actual use. First, outdoor voice interaction presents significant challenges, mainly due to concerns about privacy exposure and unstable speech recognition in noisy environments. He suggested that the system include a frequently asked questions module, similar to those used in e-commerce customer service platforms, to help users quickly access relevant information in specific situations. Second, and what he considered the most critical issue, is that although the model can recognize certain objects such as stairs, it often lacks follow-up action guidance. As a result, users are unable to safely complete tasks based solely on the provided descriptions.
Through the user study, we find that participants’ assistive needs are primarily focused on basic self-care and environmental adaptability. These include tasks such as meal preparation, medication management, personal hygiene, and mobility. Additionally, there is a strong demand for object localization and environmental awareness capabilities. From the experimental results, while most users express satisfaction with the affordability and usability of the Zhipu system, and it is able to meet some of their needs, the model still lacks sufficient accuracy and stability in its outputs.
@@
Challenge 1: Limited Performance of VideoLLMs. Although the current VideoLLMs have shown strong capabilities in real-time video understanding, user feedback, and benchmark evaluations have revealed several key issues that need to be addressed: 1. Interaction quality defects in noisy environments In complex auditory scenarios, the reliability of voice interaction of VideoLLMs is significantly reduced. Specifically, when the user is in a noisy environment, the model has difficulty in effectively extracting the user’s audio input. User feedback shows that noise interference mainly comes from two aspects: environmental background sound (such as traffic noise and crowd conversation) and clothing friction caused by user body movements (especially jackets close to mobile phones). 2. Insufficient adaptability to low-light environments: The current VideoLLMs are significantly limited in dim light or night scenes. Its visual recognition ability will be perceptibly reduced. For example, user feedback pointed out that in night street walking scenes, models such as Zhipu have difficulty accurately identifying surrounding objects and location information. 3. Fine-grained object recognition defects: VideoLLms lack the ability to distinguish objects with similar visual features. For example, the misjudgment rate of different cup types is high, and it is difficult to capture and describe subtle texture differences 4.Object description credibility issues: User study feedback that the VideoLLMs have a “hallucination” phenomenon. For example, Zhipu describes objects that do not exist in the environment in the test case. 5.Lack of consistency in language interaction: The VideoLLMs have the problem of out-of-control language switching. For example, when users query in English, Zhipu sometimes will respond in Chinese. The above problems expose the defects of the VideoLLMs in three dimensions: environmental adaptability, visual cognition accuracy, and interaction certainty. To achieve reliable deployment in real-world scenarios, key breakthroughs need to be made in VideoLLMs.
@@
Challenge 2: Lack of Capability Proactive Perception Capability. Current VideoLLMs generally adopt a passive response mechanism; that is, the model response is triggered only when the user initiates a query. That means the model cannot proactively anticipate dynamic environmental risks (such as sudden obstacles), resulting in its inability to issue warnings to users before danger occurs. To overcome this limitation, we propose SafeVid to integrate proactive environmental perception capabilities into model training. For the details of our method, please refer to the Section 5.4.
@@
Challenge 3: Limited Usability in Practice. In our experiments, interactions with VideoLLMs were conducted via smartphones using their official apps, which were worn around the neck. However, since the existing VideoLLMs are not optimized for the special needs of visually impaired users, the following difficulties are exposed in actual use: 1. Hardware interaction design flaws: The user reports that the significant sense of weight caused by the long-term wearing of the device (smartphone) affects the willingness to continue using it. To address this, a potential improvement could be to explore VideoLLMs integration solutions with lightweight wearable devices such as smart glasses; 2. Lack of privacy protection mechanism: Participants have psychological pressure to be eavesdropped when speaking personal information in public places. To address this, future research could focus on developing non-voice vibration/tactile feedback interaction channels to protect user privacy; 3. Lack Intelligent customer service: Users suggested that VideoLLMs could adopt interaction patterns from intelligent customer service systems. For example, offering proactive suggestions or providing answers to commonly asked questions could streamline user interactions, enabling them to access information more efficiently and significantly improving both the speed of interaction and the overall user experience.
@@
Challenge 4: Security Issues of VideoLLMs. Previous studies have shown that LLMs and VLMs generally have security vulnerabilities and are susceptible to backdoor attacks [36, 24] and jailbreak attacks [33, 12, 38, 19]. These vulnerabilities may pose potential risks to visually impaired individuals using such systems. Although security is not the primary focus of this study, it remains a critical concern. To explore this issue, we conduct a preliminary security evaluation by performing a jailbreak attack on the Zhipu model using the FC-Attack [38]. We find that the model could be successfully jailbroken [^4] The vulnerability has been promptly reported to the service provider for remediation.. Therefore, we argue that before VideoLLMs are widely deployed in assistive applications for visually impaired users, their security risks must be thoroughly studied and effectively addressed.
@@
As described in Section 3.2, we construct a video-text paired dataset, SafeVid, to capture potential risk levels in various environments. Based on this dataset, we modify the VITA-1.5 framework to enable the model to proactively analyze video content through timed polling and periodically generate descriptions of complicated environmental hazards. The model’s performance is evaluated using Accuracy, BLEU, and ROUGE-L.
As shown in Table 5, the original VITA-1.5 achieves an accuracy of only 26% on the environmental risk recognition task. After task-specific fine-tuning, the model’s accuracy improves to 62.24% , with a BLEU score of 22.03 and a ROUGE-L score of 36.95. This implies that fine-tuning with high high-quality dataset could enhance the VideoLLM’s proactive reminder ability significantly, which could serve as a valuable future research direction.
Table 5: Performance of fine-tuned and vanilla VITA-1.5
@@
VITA-1.5-Vanilla 26.00% 1.63 19.04
VITA-1.5-Fine-Tuned 62.24% 22.03 36.95
@@ Conclusion
This work is the first to focus on leveraging real-time interactive VideoLLMs for daily assistance to individuals with visual impairments. Building on previous research and insights from visually impaired volunteers, we construct an evaluation dataset named VisAssistDaily, and systematically evaluate three SOTA real-time VideoLLMs. Experimental results show that GPT-4o achieves the highest task success rate, while VITA-1.5 performs the worst. We then recruit eight volunteers to conduct real-world testing using the Zhipu VideoLLMs. Overall, participants express a high level of satisfaction with the system while also providing valuable feedback and highlighting several challenges. Based on these findings, we summarize key challenges currently faced by VideoLLMs in assistive applications for the visually impaired. To equip VideoLLMs with proactive risk awareness, we build a dedicated dataset (SafeVid) and fine-tune VITA-1.5, boosting risk detection accuracy to 62.24% and demonstrating its assistive potential.

### v2

@@ Results
Based on our constructed benchmark dataset VisAssistDaily, we evaluate three models: GPT-4o, Zhipu, and VITA-1.5. The results are averaged from tests conducted by three individuals wearing blindfolds to simulate visual impairments (as shown Figure˜2).
Table˜5 presents the TSR across the three models. GPT-4o performs the best in most tasks, especially within the “Home Life Tasks” and “Basic Skills”, where it achieves 100% TSR on multiple tasks. Further, the gap between its TSR in English and Chinese is minimal, indicating balanced multilingual capabilities. Zhipu performs slightly below GPT-4o but still shows strong results in both task sets, achieving an average TSR of 55.22% and 73.48%, respectively. However, its performance varies significantly across tasks. For example, it records a low success rate of only 33.33% in “Road Walking (EN)” and just 16.67% in “Cane Techniques (EN)”, indicating challenges in handling certain task-specific scenarios. VITA-1.5 shows the weakest overall performance, with a TSR of only 53.30% in English and 61.18% in Chinese. The model demonstrates inconsistent behavior across languages for different tasks. For instance, it achieves 100% TSR in “Housework (CH)”, outperforming its 65.00% in “Housework (EN)”, but only 33.33% in “Reaching Destination (EN)”, reflecting unstable multilingual task performance.
Table˜7 shows the models’ performance across various evaluation metrics. Overall, GPT-4o ranks the first, achieving the highest TSR and LC, with the lowest ARL in English. It also records the lowest APC in CH tasks.
[…]
Zhipu ranks the second, with a solid average TSR of 73.48% under CH input and moderate ARL (2.66) and APC (3.52). However, it struggles to maintain consistent language output, often responding in Chinese to EN prompts. VITA-1.5 performs the worst overall across all models.
@@
In error case 1, GPT-4o demonstrates spatial reasoning limitations, mislocating a cane positioned on the right side of the room by describing it as being on the left and near a distant black chair. This reflects confusion between object co-occurrence and actual spatial proximity, as well as unreliable left-right orientation. In error case 2, Zhipu incorrectly localizes the user at the top of a staircase despite clear contextual cues indicating the bottom position. The model further described a flat hallway ahead, omitting the presence of stairs. This reflects a failure in spatial grounding and may result in safety risks in assistive navigation contexts.
Figure 6: Example of error cases.
@@
Participants. We recruit 8 volunteers through social media platforms, including RedNotes and Weibo, to participate in the user study in either an online or in-person manner. Each participant is compensated at an hourly rate of 15 dollars. As shown in Table˜6, the participants are in a diverse range of ages (from 21 to 90). Among them, U1 to U5 are individuals who are severely visually impaired and are almost completely blind, while U6 to U8 are sighted individuals who wear blindfolds during study to simulate visual impairment.
@@
Task. The evaluation tasks are divided into two categories: closed-world tasks and open-world tasks (user-defined goals). Closed-world tasks include: (1) Finding a specific item in a room; (2) Identifying a book and helping to determine its title and table of contents; (3) Going upstairs, with the model guiding users by describing step height and assisting with navigation; (4) Entering a room and having the model describe the surroundings; (5) Walking outdoors, with the model providing feedback on road conditions. Open-world tasks include finding multiple similar objects, crossing the street, finding medicine bottles, judging road conditions and crossing traffic lights, finding tools in a workspace, assisting with learning, experiencing recreational activities, and expanding the scope of learning.
Evaluation Metrics. We use only TSR as the main metric, since some participants are not comfortable with recording video, making it difficult to collect other measurements. In addition, we use a Satisfaction score to capture overall user experience, which is rated by each participant after completing the tasks. The scoring criteria are detailed in Table˜8.
@@
Performance Analysis. As shown in Table˜6, we find that participants achieve high TSR in both closed-world and open-world settings. Notably, users U2, U6, U7, and U8 achieve a closed-world TSR of 100.00%. Besides, users U5, U4, and U3 achieve closed-world TSRs of 80.00%, 80.00%, and 75.00%, respectively. In the open-world settings, users U5, U6, and U7 achieve a TSR of 100.00%. Also, users U4, U8, and U1 attain TSRs of 83.00%, 80.00%, and 75.00%, respectively. These results suggest that Zhipu performs well in most scenarios, demonstrating high TSR across a broad range of tasks. However, a notable exception is observed with user U1 and U3, who have 60.00% and 50% closed-world TSR. This lower performance is likely due to Zhipu’s limitations in distinguishing between similar objects. Despite these exceptions, all users acknowledge the capabilities of Zhipu and are generally satisfied with the model, as demonstrated by their satisfaction scores are all above $6$. This indicates that, overall, users find the model capable of handling complex tasks to a reasonable competence.
Feedback from Participants
We organize participants’ feedback into four themes shown below.
Everyday Use Cases: Self-care and Environmental Adaptation. Participants’ assistive needs concentrate on activities of daily living, including meal preparation (U1), medication management (U1), personal hygiene (U1), and mobility (U2, U3, U5). Beyond self-care, participants also seek support for recreational activities (U5) and educational tasks (U4, U7). Across these contexts, two capabilities consistently surface as high-priority: object localization (U1–U6) and broader environmental awareness (U3, U4, U6, U8).
Wayfinding and Spatial Understanding. When seeking technological assistance for navigation, participants emphasize tasks such as finding directions (U1), recognizing rooms and locations (U3, U5), crossing streets (U1, U2), and interpreting traffic signals (U2, U3). These reports underscore object localization as a prerequisite for safe action in both indoor and outdoor spaces, and highlight the need for detailed, situation-aware scene descriptions beyond coarse labels.
Failure Modes and Interaction Pain Points. Field tests reveal inconsistent performance across scenarios. Errors are most prominent in crowded or visually complex settings, with inaccurate object or environment descriptions (U4, U6, U8). Participants also encounter imprecise spatial language (U4, U6), incorrect or incomplete object recognition (U4, U6, U8), insufficient descriptive detail for decision making (U4, U8), and difficulty translating visual information into actionable steps (U4, U6). Additional issues include occasional incorrect confirmation responses (U4), and challenges with voice interaction in public—privacy concerns and unstable automatic speech recognition (ASR) in noisy environments (U4, U8). Beyond perception and interaction, users mention slow interaction speed (U1, U8), difficulty following long or complex instructions (U7, U8), and the need for persistent reminders or guidance (U8). Figure fig.˜7 summarizes the prevalence of these issues across participants.
Deep-dive Interview of U4. In a longer follow-up interview, U4 emphasizes two concerns frequently encountered in outdoor use. First, speech-based interaction is often impractical due to privacy exposure and unreliable recognition under noise; U4 suggests a Frequency-Answer-Question (FAQ-like module for quick access to situation-specific information without extensive voice input. Second, and the most critical concern is that recognition alone (e.g., detecting a staircase) rarely suffices; users need follow-up, step-by-step guidance to complete tasks safely.
Design Implications. Taken together, these findings point to several design opportunities: (1) privacy-aware, low-audibility interaction modes (earcons/haptics) and robust ASR for noisy environments; (2) fast access to context-sensitive “FAQ” prompts or templates; (3) action-oriented outputs that translate recognition into procedural steps with safety checks; (4) clearer and more calibrated spatial descriptions (with consistent frames of reference and distance cues); (5) progressive disclosure and user-controlled pacing/summarization for long instructions; and (6) persistent guidance and reminders for safety-critical tasks.
Figure 7: Summary of common failure modes and interaction pain points reported by participants.
@@
Challenge 1: Limited Performance of VideoLLMs. Although the current VideoLLMs have shown strong capabilities in real-time video understanding, user feedback and benchmark evaluations have revealed several key issues that need to be addressed: 1. Interaction quality defects in noisy environments In complex auditory scenarios, the reliability of the voice interaction of VideoLLMs is significantly reduced. Specifically, when the user is in a noisy environment, the model has difficulty in effectively extracting the user’s audio input. User feedback shows that noise interference mainly comes from two aspects: environmental background sound (such as traffic noise and crowd conversation) and clothing friction caused by user body movements (especially jackets close to mobile phones). 2. Insufficient adaptability to low-light environments: The current VideoLLMs are significantly limited in dim light or night scenes. Its visual recognition ability will be perceptibly reduced. For example, user feedback pointed out that in night street walking scenes, models such as Zhipu have difficulty accurately identifying surrounding objects and location information. 3. Fine-grained object recognition defects: VideoLLMs lack the ability to distinguish objects with similar visual features. For example, the misjudgment rate of different cup types is high, and it is difficult to capture and describe subtle texture differences (based on our initial trial). 4. Object description credibility issues: User study feedback that the VideoLLMs have a “hallucination” phenomenon. For example, Zhipu describes objects that do not exist in the environment in the test case. 5. Lack of consistency in language interaction: The VideoLLMs have the problem of out-of-control language switching. For example, when users query in English, Zhipu sometimes will respond in Chinese. The above problems expose the defects of the VideoLLMs in three dimensions: environmental adaptability, visual cognition accuracy, and interaction certainty.
In general, although VideoLLMs have shown their ability in handling these tasks, more efforts have to be made to achieve reliable deployment in real-world scenarios.
@@
Challenge 2: Lack of Proactive Perception Capability. Current VideoLLMs generally adopt a passive response mechanism; that is, the model response is triggered only when the user initiates a query. That means the model cannot proactively anticipate dynamic environmental risks (such as sudden obstacles), resulting in its inability to issue warnings to users before danger occurs. To overcome this limitation, we propose SafeVid to integrate proactive environmental perception capabilities into model training. For the details of our method, please refer to the Section˜6.4.
@@
Challenge 3: Limited Usability in Practice. In our experiments, interactions with VideoLLMs are conducted using smartphones that are worn around the neck and run the official apps. However, since the existing VideoLLMs are not optimized for the special needs of visually impaired users, the following difficulties are exposed in actual use: 1. Hardware interaction design flaws: The user reports that the significant sense of weight caused by the long-term wearing of the device (smartphone) affects the willingness to continue using it. To address this, a potential improvement could be to explore VideoLLMs integration solutions with lightweight wearable devices such as smart glasses; 2. Lack of privacy protection mechanism: Participants have psychological pressure to be eavesdropped when speaking personal information in public places. To address this, future research could focus on developing non-voice vibration/tactile feedback interaction channels to protect user privacy; 3. Lack intelligent customer service: Users suggest that VideoLLMs can adopt interaction patterns from intelligent customer service systems. For example, offering proactive suggestions or providing answers to commonly ask questions can streamline user interactions, enabling them to access information more efficiently and significantly improving both the speed of interaction and the overall user experience.
@@
Challenge 4: Security Issues of VideoLLMs. Previous studies show that LLMs and VLMs generally have security vulnerabilities and are susceptible to backdoor attacks [45, 30] and jailbreak attacks [31, 42, 13, 47, 23]. These vulnerabilities may pose potential risks to visually impaired individuals using such systems. Although security is not the primary focus of this study, it remains a critical concern. To explore this issue, we conduct a preliminary security evaluation by performing a jailbreak attack on the Zhipu model using the FC-Attack [47]. We find that the model could be successfully jailbroken [^5] The vulnerability has been promptly reported to the service provider for remediation.. Therefore, we argue that before VideoLLMs are widely deployed in assistive applications for visually impaired users, their security risks must be thoroughly studied and effectively addressed.
@@
As described in Section˜4.2, we construct a video-text paired dataset, SafeVid, to capture potential risk levels in various environments. Based on this dataset, we modify the VITA-1.5 framework to enable the model to proactively analyze video content through timed polling and periodically generate descriptions of complicated environmental hazards. The model’s performance is evaluated using Accuracy, BLEU[27], and ROUGE-L[22]. We define accuracy as the fraction of examples for which the model’s predictions, obtained before and after fine-tuning, coincide with the ground truth labels.
As shown in  Table˜9, the original VITA-1.5 achieves an accuracy of only 25.00% on the environmental risk recognition task. After task-specific fine-tuning, the model’s accuracy improves to 76.00%, a nearly threefold improvement. Also, the fine-tuned model yields a BLEU score of 31.41 and a ROUGE-L score of 40.06, where higher values on both metrics indicate better performance. From  fig.˜8, the fine-tuned model is noticeably better at recognizing the primary risk: accuracy improves for No Risk (from 44% to 57%), Low Risk (from 20% to 68%), and Moderate Risk (from 29% to 100%). The confusion matrix indicates that the previous conservative tendency to downgrade samples has largely been corrected; remaining errors are mostly adjacent confusions between Low Risk and Moderate Risk. As a result, the model more reliably separates low from medium risk, reducing safety risks from missed or incorrect classifications. This implies that fine-tuning with high-quality dataset could enhance the VideoLLM’s proactive reminder ability significantly, which could serve as a valuable future research direction.
Table 9: Performance of vanilla and fine-tuned VITA-1.5.
@@
VITA-1.5-Vanilla 25.00% 2.76 18.35
VITA-1.5-Fine-Tuned 76.00% 31.41 40.06
@@ Conclusion
This work is the first to focus on leveraging real-time interactive VideoLLMs for daily assistance to individuals with visual impairments. Building on previous research and insights from visually impaired volunteers, we construct an evaluation dataset named VisAssistDaily, and systematically evaluate three SOTA real-time VideoLLMs. Experimental results show that GPT-4o achieves the highest task success rate, while VITA-1.5 performs the worst. We then recruit eight volunteers to conduct real-world testing using the Zhipu VideoLLMs. Overall, participants express a high level of satisfaction with the system while also providing valuable feedback and highlighting several challenges. Based on these findings, we summarize key challenges currently faced by VideoLLMs in assistive applications for the visually impaired. To equip VideoLLMs with proactive risk awareness, we build a dedicated dataset (SafeVid) and fine-tune VITA-1.5, boosting risk detection accuracy to 76.00% and demonstrating its assistive potential.